OpenStudy (anonymous):

Which point lies on the line described by the equation below? y + 4 = 4(x - 3)

terenzreignz (terenzreignz):

I'm assuming you have choices?

OpenStudy (anonymous):

a(2, 9) b(3, -4) c(1, -11) d(1, 11) e(0, 0) f(-2, 3)

terenzreignz (terenzreignz):

That's better... Now here's how to do it, check each point (there's no other way, unfortunately).... for example, a. (2 , 9) Now, replace x with 2, and y with 9 And see if things "make sense" 9 + 4 = 4(2 - 3) 13 = 4(-1) 13 = -4 And it doesn't make sense (it should lead to a reasonable equation, like 6 = 6) So a(2,9) is not the answer :P Go now, try the others...

OpenStudy (anonymous):

Thanks again (:

terenzreignz (terenzreignz):

No problem :)